1. Represent the following sets in tabular or listing form:


(i) A={x:x^2 - 3x + 2 = 0}


(ii) B= { x:x is an integer and 1 < x < 7}



1
Expert's answer
2021-11-04T08:08:50-0400

i)

"x^2 - 3x + 2 = 0"

"x=\frac{3\pm \sqrt{9-8}}{2}"

"x=\{1,2\}"


ii)

"x=\{2,3,4,5,6\}"